Suppose a 250. mL flask is filled with 0.10 mol of Cl2 and 1.4 mol of HCl. The following reaction becomes possible:

H2(g)+Cl2(g)=2HCl(g)

The equilibrium constant for this reaction is 0.414 at the temperature of the flask.

Calculate the equilibrium molarity of . Round your answer to two decimal places.
Answer #1

you of want H₂ cg to at find the molar concentration equilibrium. molar concentration of de a 0.10 0.40M 0.250 molar concentrgo H₂ & clog o 0.40m 2 Ha y 5.60m C + x + x - 2x E (x) (0.40+x) (5.60-2x) ke- [HAJ? [de] Hg 0,4 14 - - (5.60-2x e (0.404) (x2I Hal = 5.60 - 2x = 5.60-9x4.042 = -ve. bo bo, it is negligible , then this value is not considered So the value of x = 2.166